### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: the 32th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 09–13, 2026; Jeju, Koreaisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Neural networks
@@ 1. Introduction
Forecasting is a notably critical problem in the time series analysis community, which aims to predict future time series based on historical time series records (wang2024deep). It has broad practical applications such as climate modeling (wu2023interpretable), traffic flow management (yin2021deep), healthcare monitoring (kaushik2020ai) and finance analytics (granger2014forecasting).
@@
rise of Transformers in various domains (DBLP:journals/corr/abs-1810-04805; khan2022transformers) has led to their wide adoption in time series forecasting (wen2022transformers; autoformer; fedformer; informer; patchtst; itransformer), leveraging the strong capabilities of capturing pairwise dependencies and extracting multi-level representations within sequential data. Similarly, CNN architectures have also proven effective by developing convolution blocks for time series (moderntcn; micn). However, despite the strong performance of Transformer-based and CNN-based models, they face significant challenges in large-scale industrial applications due to their relatively high computational demands, especially in latency-sensitive scenarios like financial prediction and healthcare monitoring (granger2014forecasting; kaushik2020ai). In contrast, simpler linear or MLP models offer greater efficiency, although with lower performance (dlinear; sparsetsf). These contrasting observations raises an intriguing question:
@@
A promising approach to addressing this question is knowledge distillation (KD(hinton2015distilling), a technique that transfers knowledge from a larger and more complex model (teacher) to a smaller and simpler one (student) while maintaining comparable performance. In this work, we pioneer cross-architecture KD in time series forecasting, with MLP as the student and other advanced architectures (e.g., Transformers and CNNs) as the teacher. However, designing such a framework is non-trivial, as it remains unclear what “knowledge” should be distilled into MLP.
To investigate this potential, we conduct a comparative analysis of prediction patterns between MLP and other time series models. Our findings reveal that MLP still excels on some data subsets despite its overall lower performance (Sec. 3.1), which highlights the value of harnessing the complementary capabilities across different architectures. To further explore the specific properties to distill, we focus on two key time series patterns: multi-scale pattern in temporal domain and multi-period pattern in frequency domain, given that they are vital in capturing the complex structures typical of many time series. (1) Multi-Scale Pattern: Real-world time series often show variations at multiple temporal scales. For example, hourly recorded traffic flow data capture changes within each day, while daily sampled data lose fine-grained details but reveal patterns related to holidays (timemixer). We observe that models that perform well on the finest scale also perform accurately on coarser scales, while MLP fails on most scales (Sec. 3.2). (2) Multi-Period Pattern: Time series often exhibit multiple periodicities. For instance, weather measurements may have both daily and yearly cycles, while electricity consumption data may show weekly and quarterly cycles (timesnet). We find that models that perform well can capture periodicities similar to those in the ground truth, but MLP fails to capture these periodicities (Sec. 3.2). Therefore, enhancing MLP requires distilling and integrating these multi-scale and multi-period patterns from teacher models.
@@ 2.1. Debate in Long-Term Time Series Forecast
The trade-off between performance and efficiency has prompted a long-standing debate between Transformer-based models and MLP in long-term time series forecasting (dlinear; lightts; sparsetsf). Informer (informer), Autoformer (autoformer), and FEDformer (fedformer) were among the leading Transformer-based methods. However, recent findings show that a simple Linear or MLP model can achieve performance comparable to these complex Transformer models across various benchmarks while offering significantly better efficiency (dlinear). This outcome has raised questions about the necessity of Transformers in time series forecasting. Following this, research has moved in two directions. One direction suggests that the issues with Transformer-based models arise from the way they are applied. For example, PatchTST (patchtst) uses patching to preserve local information, and iTransformer (itransformer) focuses on capturing multivariate correlations. These approaches surpass the simple one-layer MLP and demonstrate that Transformers can still deliver strong results in time series forecasting if applied effectively. Meanwhile, CNN-based models have also shown strong performance similar to Transformer-based models. TimesNet (timesnet) transforms 1D time series into 2D variations and applies 2D CNN kernels, MICN (micn) adopt multi-scale convolution structures to capture local features and global correlations, and ModernTCN (moderntcn) proposes a framework with much larger receptive fields than prior CNN-based structures. Nevertheless, these powerful CNN-based models also face efficiency issues, which further broadens the scope of the debate between performance and efficiency. The other direction focuses on developing lightweight MLP, such as N-BEATS (nbeats), N-hits (nhits), LightTS (lightts), FreTS (frets), TSMixer (tsmixer), TiDE (tide), which offer improved efficiency. However, these models typically only match, rather than surpass, the performance of state-of-the-art Transformer-based methods and CNN-based methods. In summary, while Transformer-based and CNN-based models generally offer better performance, simple MLP is more efficient. Therefore, we managed to combine the performance of Transformer-based and CNN-based models with MLP to produce a powerful and efficient model by cross-architecture KD.
@@ 2.2. Knolwedge Distillation on Time Series
Knowledge distillation (KD(hinton2015distilling) transfers knowledge from a larger, more complex model (teacher) to a smaller, simpler model (student) while maintaining comparable performance. By aligning the output distributions of teacher and student models, KD provides richer training signals than hard labels alone, enabling the student to capture subtle patterns that the teacher has learned. In the context of time series, CAKD (xu2022contrastive) introduces a two-stage distillation scheme that distills features using adversarial and contrastive learning and performs prediction-level distillation. LightTS (campos2023lightts) designs a KD framework specifically for cases where the teacher is an ensemble classifier in time series classification tasks, which limits its applicability to teachers with other architectures. Both of these works do not incorporate time series-specific designs. In contrast, our framework emphasizes extracting essential time series patterns, including multi-scale and multi-period patterns, enabling more effective knowledge distillation. Additionally, we are the first to explore cross-architecture KD.
@@ 3.1. What Motivates Distillation?
An intuitive strategy of KD for MLP is to align predictions with teachers, but this faces limitations. First, it risks overfitting noise in the teacher’s predictions, leading to less stable knowledge (chen2017learning; takamoto2020efficient; gajbhiye2021knowledge). Second, MLP may struggle to replicate the complex patterns such as seasonality, trends, and periodicity in teacher predictions directly due to their limited capacity. Third, this approach overlooks valuable knowledge from intermediate features of teacher model. Therefore, the specific knowledge to distill into MLP requires further exploration.
@@ 3.2. What Should We Distill?
To investigate which complementary time series patterns to distill into MLP, we analyze prediction patterns in both temporal and frequency domains, considering real-world variations across temporal scales (timemixer) and periodicities (timesnet). Thus, we conduct further analysis of these two patterns by presenting illustrative cases. The implementation is detailed in Appendix E. (1) Multi-Scale Pattern: By downsampling predictions $\mathbf{\hat{Y}}$ using convolutional operations, we obtain their multi-scale representations. Figure 5 shows the analysis of multi-scale temporal patterns by downsampling the time series (Scale 0) to coarser scales (Scales 1–3). Models excelling at Scale 0 generally perform well across all scales. The teacher models align closely with the ground truth at Scale 3, capturing underlying trends effectively. In contrast, MLP significantly deviates, showing its limitations in handling multi-scale patterns. (2) Multi-Period Pattern: Periodicities in time series are visible in the frequency domain by transforming time series into spectrograms, with the x-axis representing frequency and periodicity calculated as the time series length $S$ divided by frequency. Figure 5 shows spectrograms of model predictions and the ground truth. Models with lower MSE capture periodicities more accurately, closely matching the ground truth at dominant frequencies. In contrast, MLP predictions show larger discrepancies, highlighting its inability to effectively capture periodic patterns.
@@ 4. Methodology
where $\theta_{s}$ is the parameter of the student; $\mathcal{L}_{sup}$ is the supervised loss (e.g., MSE) between predictions and ground truth; $\mathcal{L}_{\mathrm{KD}}^{\mathbf{Y}}$ and $\mathcal{L}_{\mathrm{KD}}^{\mathbf{H}}$ are the distillation loss terms that encourage student model to learn knowledge from teacher on both prediction level (hinton2015distilling) and feature level (romero2014fitnets). Unlike conventional approaches that emphasize matching model predictions, TimeDistill integrates key time-series patterns including multi-scale and multi-period knowledge. The overall framework of TimeDistill is shown in Figure 6.
@@ 4.2. Multi-Period Distillation
To reduce the influence of minor frequencies and avoid noise introduced by less meaningful frequencies (timesnet; fedformer), we propose a distribution-based matching scheme. We use softmax function with a colder temperature to highlight the most significant frequencies:
@@
where $\mathbf{Q}_{x}^{\mathbf{Y}}\in\mathbb{R}^{\frac{S}{2}\times C}$ and $\tau$ is a temperature parameter that controls the sharpness of the distribution. We set $\tau=0.5$ by default. The period distribution $\mathbf{Q}_{x}^{\mathbf{Y}}$ represents the multi-period pattern in the prediction time series, which we want the MLP to learn from the teacher. We use KL divergence to match these distributions (hinton2015distilling). We define multi-period distillation loss at the prediction level as:
@@ 4.3. Overall Optimization and Theoretical Analysis
Theoretical Interpretations. We provide a theoretical understanding of multi-scale and multi-period distillation loss from a novel data augmentation perspective. We further show that the proposed distillation loss can be interpreted as training with augmented samples derived from a special mixup (mixup) strategy. The distillation process augments data by blending ground truth with teacher predictions, analogous to label smoothing in classification, and provides several benefits for time series forecasting: (1) Enhanced Generalization: It enhances generalization by exposing the student model to richer supervision signals from augmented samples, thus mitigating overfitting, especially with limited or noisy data. (2) Explicit Integration of Patterns: The augmented supervision signals explicitly incorporate patterns across multiple scales and periods, offering insights that are not immediately evident in the raw ground truth. (3) Stabilized Training Dynamics: The blending of targets softens the supervision signals, which diminishes the model’s sensitivity to noise and leads to more stable training phases. This will in turn support smoother optimization dynamics and fosters improved convergence. For clarity, our discussion is centered at the prediction level. We present the following theorem:
@@
Let $(x,y)$ denote original input data pairs and $(x,y^{t})$ represent corresponding teacher data pairs. Consider a data augmentation function $\mathcal{A}(\cdot)$ applied to $(x,y)$, generating augmented samples $(x^{\prime},y^{\prime})$. Define the training loss on these augmented samples as $\mathcal{L}_{aug}=\sum_{(x^{\prime},y^{\prime})\in\mathcal{A}(x,y)}|f_{s}(x^{\prime})-y^{\prime}|^{2}$. Then, the following inequality holds: $\mathcal{L}_{sup}+\eta\mathcal{L}_{scale}\geq\mathcal{L}_{aug},$ when $\mathcal{A}(\cdot)$ is instantiated as a mixup function (mixup) that interpolates between the original input data $(x,y)$ and teacher data $(x,y^{t})$ with a mixing coefficient $\lambda=\frac{1}{1+\eta}$, i.e. $y^{\prime}=\lambda y+(1-\lambda)y^{t}$.
@@ 5. Experiment
(Ours) ((itransformer)) ((moderntcn)) ((timemixer)) ((patchtst)) ((micn)) ((fedformer)) ((timesnet)) ((autoformer))
@@ 5.1. Experimental Setup
Datasets and Baselines. We run experiments to evaluate the performance and efficiency of TimeDistill on 8 widely used benchmarks: Electricity (ECL), the ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2), Solar, Traffic, and Weather, following (itransformer; timemixer; moderntcn). To examine the effectiveness of our method across diverse tasks, we compare TimeDistill with 8 baseline models that cover a range of architectures. Specifically, we use Transformer-based models: iTransformer (itransformer), PatchTST (patchtst), FEDformer (fedformer), Autoformer (autoformer); CNN-based models: ModernTCN (moderntcn), MICN (micn), TimesNet (timesnet); and an MLPs-based model: TimeMixer (timemixer). More details can be found in Appendix B.
@@ 5.3. Versatility of TimeDistill
Different Teachers. We use ModernTCN as the teacher model for our main results. In Table 2, we present results using other teacher models, such as iTransformer, TimeMixer, and PatchTST. Results with additional teachers are provided in Appendix F. We observe that TimeDistill can effectively learn from various teachers, improving MLP by up to 13.87%. Furthermore, TimeDistill achieves significant performance improvements over the teachers themselves, with gains of up to 21.41%. We hypothesize two key reasons for these improvements. First, the student MLP model already demonstrates strong learning capabilities; for instance, on the Solar dataset, MLP outperforms most teachers. Second, the multi-scale and multi-period KD approach delivers both temporal and frequency knowledge from teachers to MLP, offering additional valuable insights. Similar findings have been reported in recent KD studies (allen2020towards; zhang2021graph; guo2023linkless), which suggest that integrating diverse views from multiple models can enhance performance.
Different Students. To evaluate whether TimeDistill can enhance the performance of other lightweight models, we select two MLPs-based models, TSMixer (tsmixer), LightTS (lightts), as students. Unlike a simple MLP, these MLPs-based models consist of multiple MLPs and operate in a channel-dependent manner. Consequently, they are generally more complex than a standard MLP, which results in reduced efficiency. Besides, we select a Linear-based model FITS (fits), which is extremely lightweight. We use ModernTCN as the teacher model, consistent with other experiments. As shown in Table 3, TimeDistill consistently improves the performance of TSMixer, LightTS, and FITS, achieving remarkable MSE reductions of 6.26%, 8.02%, and 3.96%, respectively. These results demonstrate that TimeDistill is highly adaptable and can effectively enhance other lightweight methods. We also explore the influence of student MLP architecture (e.g. layer number and hidden dimension) on TimeDistill in Appendix G.

### v3

@@ Abstract.
* Corresponding author.
copyright: acmlicensedjournalyear: 2026copyright: ccconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 09–13, 2026; Jeju Island, Republic of Koreabooktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’26), August 09–13, 2026, Jeju Island, Republic of Koreadoi: 10.1145/3770854.3780218isbn: 979-8-4007-2258-5/2026/08ccs: Information systems Temporal dataccs: Mathematics of computing Time series analysisccs: Computing methodologies Neural networks
@@ 1. Introduction
Forecasting is a notably critical problem in the time series analysis community, which aims to predict future time series based on historical time series records (Wang et al., 2024b). It has broad practical applications such as climate modeling (Wu et al., 2023), traffic flow management (Yin et al., 2021), healthcare monitoring (Kaushik et al., 2020) and finance analytics (Granger and Newbold, 2014).
@@
rise of Transformers in various domains (Devlin et al., 2018; Khan et al., 2022) has led to their wide adoption in time series forecasting (Wen et al., 2022; Wu et al., 2021; Zhou et al., 2022, 2021; Nie et al., 2023; Liu et al., 2024), leveraging the strong capabilities of capturing pairwise dependencies and extracting multi-level representations within sequential data. Similarly, CNN architectures have also proven effective by developing convolution blocks for time series (Luo and Wang, 2024; Wang et al., 2023). However, despite the strong performance of Transformer-based and CNN-based models, they face significant challenges in large-scale industrial applications due to their relatively high computational demands, especially in latency-sensitive scenarios like financial prediction and healthcare monitoring (Granger and Newbold, 2014; Kaushik et al., 2020). In contrast, simpler linear or MLP models offer greater efficiency, although with lower performance (Zeng et al., 2023; Lin et al., 2024). These contrasting observations raises an intriguing question:
@@
A promising approach to addressing this question is knowledge distillation (KD(Hinton, 2015), a technique that transfers knowledge from a larger and more complex model (teacher) to a smaller and simpler one (student) while maintaining comparable performance. In this work, we pioneer cross-architecture KD in time series forecasting, with MLP as the student and other advanced architectures (e.g., Transformers and CNNs) as the teacher. However, designing such a framework is non-trivial, as it remains unclear what “knowledge” should be distilled into MLP.
To investigate this potential, we conduct a comparative analysis of prediction patterns between MLP and other time series models. Our findings reveal that MLP still excels on some data subsets despite its overall lower performance (Sec. 3.1), which highlights the value of harnessing the complementary capabilities across different architectures. To further explore the specific properties to distill, we focus on two key time series patterns: multi-scale pattern in temporal domain and multi-period pattern in frequency domain, given that they are vital in capturing the complex structures typical of many time series. (1) Multi-Scale Pattern: Real-world time series often show variations at multiple temporal scales. For example, hourly recorded traffic flow data capture changes within each day, while daily sampled data lose fine-grained details but reveal patterns related to holidays (Wang et al., 2024a). We observe that models that perform well on the finest scale also perform accurately on coarser scales, while MLP fails on most scales (Sec. 3.2). (2) Multi-Period Pattern: Time series often exhibit multiple periodicities. For instance, weather measurements may have both daily and yearly cycles, while electricity consumption data may show weekly and quarterly cycles (Wu et al., 2022). We find that models that perform well can capture periodicities similar to those in the ground truth, but MLP fails to capture these periodicities (Sec. 3.2). Therefore, enhancing MLP requires distilling and integrating these multi-scale and multi-period patterns from teacher models.
@@ 2.1. Debate in Long-Term Time Series Forecast
The trade-off between performance and efficiency has prompted a long-standing debate between Transformer-based models and MLP in long-term time series forecasting (Zeng et al., 2023; Zhang et al., 2022; Lin et al., 2024). Informer (Zhou et al., 2021), Autoformer (Wu et al., 2021), and FEDformer (Zhou et al., 2022) were among the leading Transformer-based methods. However, recent findings show that a simple Linear or MLP model can achieve performance comparable to these complex Transformer models across various benchmarks while offering significantly better efficiency (Zeng et al., 2023). This outcome has raised questions about the necessity of Transformers in time series forecasting. Following this, research has moved in two directions. One direction suggests that the issues with Transformer-based models arise from the way they are applied. For example, PatchTST (Nie et al., 2023) uses patching to preserve local information, and iTransformer (Liu et al., 2024) focuses on capturing multivariate correlations. These approaches surpass the simple one-layer MLP and demonstrate that Transformers can still deliver strong results in time series forecasting if applied effectively. Meanwhile, CNN-based models have also shown strong performance similar to Transformer-based models. TimesNet (Wu et al., 2022) transforms 1D time series into 2D variations and applies 2D CNN kernels, MICN (Wang et al., 2023) adopt multi-scale convolution structures to capture local features and global correlations, and ModernTCN (Luo and Wang, 2024) proposes a framework with much larger receptive fields than prior CNN-based structures. Nevertheless, these powerful CNN-based models also face efficiency issues, which further broadens the scope of the debate between performance and efficiency. The other direction focuses on developing lightweight MLP, such as N-BEATS (Oreshkin et al., 2020), N-hits (Challu et al., 2023), LightTS (Zhang et al., 2022), FreTS (Yi et al., 2023), TSMixer (Chen et al., 2023), TiDE (Das et al., 2023), which offer improved efficiency. However, these models typically only match, rather than surpass, the performance of state-of-the-art Transformer-based methods and CNN-based methods. In summary, while Transformer-based and CNN-based models generally offer better performance, simple MLP is more efficient. Therefore, we managed to combine the performance of Transformer-based and CNN-based models with MLP to produce a powerful and efficient model by cross-architecture KD.
@@ 2.2. Knolwedge Distillation on Time Series
Knowledge distillation (KD(Hinton, 2015) transfers knowledge from a larger, more complex model (teacher) to a smaller, simpler model (student) while maintaining comparable performance. By aligning the output distributions of teacher and student models, KD provides richer training signals than hard labels alone, enabling the student to capture subtle patterns that the teacher has learned. In the context of time series, CAKD (Xu et al., 2022) introduces a two-stage distillation scheme that distills features using adversarial and contrastive learning and performs prediction-level distillation. LightTS (Campos et al., 2023) designs a KD framework specifically for cases where the teacher is an ensemble classifier in time series classification tasks, which limits its applicability to teachers with other architectures. Both of these works do not incorporate time series-specific designs. In contrast, our framework emphasizes extracting essential time series patterns, including multi-scale and multi-period patterns, enabling more effective knowledge distillation. Additionally, we are the first to explore cross-architecture KD.
@@ 3.1. What Motivates Distillation?
An intuitive strategy of KD for MLP is to align predictions with teachers, but this faces limitations. First, it risks overfitting noise in the teacher’s predictions, leading to less stable knowledge (Chen et al., 2017; Takamoto et al., 2020; Gajbhiye et al., 2021). Second, MLP may struggle to replicate the complex patterns such as seasonality, trends, and periodicity in teacher predictions directly due to their limited capacity. Third, this approach overlooks valuable knowledge from intermediate features of teacher model. Therefore, the specific knowledge to distill into MLP requires further exploration.
@@ 3.2. What Should We Distill?
To investigate which complementary time series patterns to distill into MLP, we analyze prediction patterns in both temporal and frequency domains, considering real-world variations across temporal scales (Wang et al., 2024a) and periodicities (Wu et al., 2022). Thus, we conduct further analysis of these two patterns by presenting illustrative cases. The implementation is detailed in Appendix E. (1) Multi-Scale Pattern: By downsampling predictions $\mathbf{\hat{Y}}$ using convolutional operations, we obtain their multi-scale representations. Figure 5 shows the analysis of multi-scale temporal patterns by downsampling the time series (Scale 0) to coarser scales (Scales 1–3). Models excelling at Scale 0 generally perform well across all scales. The teacher models align closely with the ground truth at Scale 3, capturing underlying trends effectively. In contrast, MLP significantly deviates, showing its limitations in handling multi-scale patterns. (2) Multi-Period Pattern: Periodicities in time series are visible in the frequency domain by transforming time series into spectrograms, with the x-axis representing frequency and periodicity calculated as the time series length $S$ divided by frequency. Figure 5 shows spectrograms of model predictions and the ground truth. Models with lower MSE capture periodicities more accurately, closely matching the ground truth at dominant frequencies. In contrast, MLP predictions show larger discrepancies, highlighting its inability to effectively capture periodic patterns.
@@ 4. Methodology
where $\theta_{s}$ is the parameter of the student; $\mathcal{L}_{sup}$ is the supervised loss (e.g., MSE) between predictions and ground truth; $\mathcal{L}_{\mathrm{KD}}^{\mathbf{Y}}$ and $\mathcal{L}_{\mathrm{KD}}^{\mathbf{H}}$ are the distillation loss terms that encourage student model to learn knowledge from teacher on both prediction level (Hinton, 2015) and feature level (Romero et al., 2014). Unlike conventional approaches that emphasize matching model predictions, TimeDistill integrates key time-series patterns including multi-scale and multi-period knowledge. The overall framework of TimeDistill is shown in Figure 6.
@@ 4.2. Multi-Period Distillation
To reduce the influence of minor frequencies and avoid noise introduced by less meaningful frequencies (Wu et al., 2022; Zhou et al., 2022), we propose a distribution-based matching scheme. We use softmax function with a colder temperature to highlight the most significant frequencies:
@@
where $\mathbf{Q}_{x}^{\mathbf{Y}}\in\mathbb{R}^{\frac{S}{2}\times C}$ and $\tau$ is a temperature parameter that controls the sharpness of the distribution. We set $\tau=0.5$ by default. The period distribution $\mathbf{Q}_{x}^{\mathbf{Y}}$ represents the multi-period pattern in the prediction time series, which we want the MLP to learn from the teacher. We use KL divergence to match these distributions (Hinton, 2015). We define multi-period distillation loss at the prediction level as:
@@ 4.3. Overall Optimization and Theoretical Analysis
Theoretical Interpretations. We provide a theoretical understanding of multi-scale and multi-period distillation loss from a novel data augmentation perspective. We further show that the proposed distillation loss can be interpreted as training with augmented samples derived from a special mixup (Zhang, 2017) strategy. The distillation process augments data by blending ground truth with teacher predictions, analogous to label smoothing in classification, and provides several benefits for time series forecasting: (1) Enhanced Generalization: It enhances generalization by exposing the student model to richer supervision signals from augmented samples, thus mitigating overfitting, especially with limited or noisy data. (2) Explicit Integration of Patterns: The augmented supervision signals explicitly incorporate patterns across multiple scales and periods, offering insights that are not immediately evident in the raw ground truth. (3) Stabilized Training Dynamics: The blending of targets softens the supervision signals, which diminishes the model’s sensitivity to noise and leads to more stable training phases. This will in turn support smoother optimization dynamics and fosters improved convergence. For clarity, our discussion is centered at the prediction level. We present the following theorem:
@@
Let $(x,y)$ denote original input data pairs and $(x,y^{t})$ represent corresponding teacher data pairs. Consider a data augmentation function $\mathcal{A}(\cdot)$ applied to $(x,y)$, generating augmented samples $(x^{\prime},y^{\prime})$. Define the training loss on these augmented samples as $\mathcal{L}_{aug}=\sum_{(x^{\prime},y^{\prime})\in\mathcal{A}(x,y)}|f_{s}(x^{\prime})-y^{\prime}|^{2}$. Then, the following inequality holds: $\mathcal{L}_{sup}+\eta\mathcal{L}_{scale}\geq\mathcal{L}_{aug},$ when $\mathcal{A}(\cdot)$ is instantiated as a mixup function (Zhang, 2017) that interpolates between the original input data $(x,y)$ and teacher data $(x,y^{t})$ with a mixing coefficient $\lambda=\frac{1}{1+\eta}$, i.e. $y^{\prime}=\lambda y+(1-\lambda)y^{t}$.
@@ 5. Experiment
(Ours) ((2024)) ((2024)) ((2024a)) ((2023)) ((2023)) ((2022)) ((2022)) ((2021))
@@ 5.1. Experimental Setup
Datasets and Baselines. We run experiments to evaluate the performance and efficiency of TimeDistill on 8 widely used benchmarks: Electricity (ECL), the ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2), Solar, Traffic, and Weather, following (Liu et al., 2024; Wang et al., 2024a; Luo and Wang, 2024). To examine the effectiveness of our method across diverse tasks, we compare TimeDistill with 8 baseline models that cover a range of architectures. Specifically, we use Transformer-based models: iTransformer (Liu et al., 2024), PatchTST (Nie et al., 2023), FEDformer (Zhou et al., 2022), Autoformer (Wu et al., 2021); CNN-based models: ModernTCN (Luo and Wang, 2024), MICN (Wang et al., 2023), TimesNet (Wu et al., 2022); and an MLPs-based model: TimeMixer (Wang et al., 2024a). More details can be found in Appendix B.
@@ 5.3. Versatility of TimeDistill
Different Teachers. We use ModernTCN as the teacher model for our main results. In Table 2, we present results using other teacher models, such as iTransformer, TimeMixer, and PatchTST. Results with additional teachers are provided in Appendix F. We observe that TimeDistill can effectively learn from various teachers, improving MLP by up to 13.87%. Furthermore, TimeDistill achieves significant performance improvements over the teachers themselves, with gains of up to 21.41%. We hypothesize two key reasons for these improvements. First, the student MLP model already demonstrates strong learning capabilities; for instance, on the Solar dataset, MLP outperforms most teachers. Second, the multi-scale and multi-period KD approach delivers both temporal and frequency knowledge from teachers to MLP, offering additional valuable insights. Similar findings have been reported in recent KD studies (Allen-Zhu and Li, 2020; Zhang et al., 2021; Guo et al., 2023), which suggest that integrating diverse views from multiple models can enhance performance.
Different Students. To evaluate whether TimeDistill can enhance the performance of other lightweight models, we select two MLPs-based models, TSMixer (Chen et al., 2023), LightTS (Zhang et al., 2022), as students. Unlike a simple MLP, these MLPs-based models consist of multiple MLPs and operate in a channel-dependent manner. Consequently, they are generally more complex than a standard MLP, which results in reduced efficiency. Besides, we select a Linear-based model FITS (Xu et al., 2023), which is extremely lightweight. We use ModernTCN as the teacher model, consistent with other experiments. As shown in Table 3, TimeDistill consistently improves the performance of TSMixer, LightTS, and FITS, achieving remarkable MSE reductions of 6.26%, 8.02%, and 3.96%, respectively. These results demonstrate that TimeDistill is highly adaptable and can effectively enhance other lightweight methods. We also explore the influence of student MLP architecture (e.g. layer number and hidden dimension) on TimeDistill in Appendix G.